40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1666     Entry time: Wed Jun 10 09:28:14 2009
Author: steve 
Type: Update 
Category: PSL 
Subject: PMC 

The PMC alarm was on this morning. It was relocked at lower HV

The FSS_RMTEMP jumped 0.5 C so The PZT was compensating for it.

 

Attachment 1: pmctemphv.jpg  80 kB  | Show | Show all
ELOG V3.1.3-